### v1

# Signatures of Jet Drift in QGP Hard Probe Observables

Abstract
Hard probe tomography of the quark-gluon plasma (QGP) in heavy ion collisions has long been a preeminent goal of the high-energy nuclear physics program. In service of this goal, the isotropic modification of jets and high-energy hadrons has been studied in great detail at the leading-power (eikonal) level, with effects originating from sub-eikonal anisotropic interactions presumed to be small. We present the first investigation of sub-eikonal, collective-flow-induced asymmetric jet broadening (jet drift) in event-by-event TeV PbPb collisions at the Large Hadron Collider using the new Anisotropic Parton Evolution (APE) computational framework. We show that jet drift imparts a sizeable enhancement of elliptic flow () and increases the mean acoplanarity for low and intermediate energy particles ( GeV). Importantly, these novel modifications to hard probe observables are shown to survive averaging over events and collision geometry. They couple to the collective flow of the medium seen by the jet and encode information about the QGP dynamics inaccessible to studies considering only isotropic, eikonal level effects.
I Introduction.
Hard probes are a cornerstone in the investigation of the novel form of nuclear matter – the quark-gluon plasma – produced in heavy ion collisions at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) (see [1] for a review). Following initial and acclaimed success [2], progress into hard probe tomography, especially recently, has been limited by the inherent complexity of the problem. The suppression of energetic particles and jets, measured by the nuclear modification factor
where is the number of binary collisions, is a robust quantity which can be reproduced by a wide variety of pQCD calculations and other models [3, 4, 5, 6, 7, 8] (see [9] for a direct multi-model comparison). Similarly, measurements of isotropic transverse momentum broadening like two-particle away-side correlations are complicated by competing background effects like energy-resolution smearing and medium response [1, 10]. In contrast, there are few, if any, background effects that can mimic the anisotropic coupling of energetic partons to preferred directions in the medium. Anisotropic modifications which go beyond the leading-power eikonal limit are, thus, more discriminating with respect the microscopic Quantum Chromodynamics (QCD) interactions in matter than their isotropic counterparts and encode unique information about the quark-gluon plasma (QGP).
In previous work [11], as part of the effort to develop the full theoretical framework of QCD dynamics in matter, the authors have shown that the momentum broadening of a hard parton has an anisotropic component, being a medium scale and the parton energy, which couples to the local collective flow velocity in a thermal QGP. On average, this results in a net deflection of the parton in the direction of : the so called “jet drift” effect. The aggregation of the locally anisotropic jet drift effect over the path length creates a global correlation between the inclusive azimuthal distribution of hard particles and the geometric anisotropy of the medium (the eccentricity ), enhancing the corresponding anisotropic flow of hard particles [12] defined as
where is the angle of the harmonic hard particle event plane [3]. We emphasize that jet drift represents an additional level of coupling to the medium geometry, distinct from the sensitivity to the path length dependence, although jet drift does nontrivially mix with these effects. However, at large parton energy suggests that modifications of hard probe observables might be small. Furthermore, one may reasonably question whether anisotropies due to the medium velocity field survive averaging over fluctuating events. Thus, it is imperative to investigate this question on an event-by-event level. We note that power-suppressed (“sub-eikonal”) corrections like jet drift can modify every part of the jet production cross section: the hard scattering, the medium matrix elements, and the hadronization process. Jet drift, which selects on locally asymmetric interactions that can only come from the medium, therefore provides a first window into effects that go beyond the well-understood, leading-power formalism.
In this work, we present the first study of event-by-event jet drift via APE (Anisotropic Partonic Evolution), a new open-source Monte Carlo parton trajectory simulator, that incorporates a host of flexible model choices to survey the variability of the effect. We demonstrate conclusively that even using the most conservative model choices the imprint of flow on hard probes is robust and survives averaging over many events on the example of TeV PbPb collisions at the Large Hadron Collider (LHC). Furthermore we show that well-motivated modifications to hadronization and medium modeling have the potential to further enhance the effect of jet drift, showing drastic modification to hard observables below hadronic of 10 GeV. Importantly, this new framework differs substantially from most other similar hard particle and jet simulators (e.g., HIJING [13], HYDJET++ [14], JEWEL [15], JETSCAPE [16], DREENA-A [17], HYBRID [18]) by tracking step-by-step average deflections to the trajectory of hard particles as they propagate through the evolving medium. In this work, we present the first serious, event-by-event simulation of flow effects from genuine power-suppressed corrections, which goes well beyond approaches which simply boost the leading-power (“eikonal”) scattering amplitudes to a moving frame [13, 19, 18].
II Methodology.
Our realistic event-by-event medium is a re-implementation of the open-source DukeQCD hybrid transport model [20], which consists of an initial condition from ENTO [21, 22], 2+1D viscous hydrodynamics and Cooper-Frye particlization [23] followed by the UrQMD hadronic afterburner [24]. Consistent with the boost-invariant medium, we treat the mid-rapidity production of hard partons as boost invariant for a given 111When discussing the power counting of the eikonal expansion, we refer to the parton energy rather than to avoid confusion with drift transverse to the jet axis. For light particle production at mid rapidity , the two are equivalent: . We employ the high-probability parameters used to tune the model from Ref. [20], and we verify that our simulations quantitatively reproduce the soft , , and flow harmonics as well as identified hadron multiplicities, mean , and fluctuations reported therein.
The seeds of jets in APE are modeled as single partons, for which we calculate the drift and energy loss. This approach effectively neglects splittings other than medium-induced soft gluon emission, under the assumption that the hard scattering, in medium interactions, and fragmentation are factorizable [2]. In many Monte Carlo simulations, a relatively small number of QGP geometries are each seeded with many hard-scattering events for computational efficiency (see, e.g., [3, 26]). In contrast, we sample only 1000 hard-scattering events per event geometry, but we simulate a much larger number of events. This lower oversampling factor retains the correlations within an event geometry as much as possible, while still saturating the sampling of the phase space.
Hard processes from TeV Pythia [27] pp collisions are sampled using an importance envelope of for a parton with momentum . Equivalently, this assigns a weight to each trajectory, allowing us to maintain controlled statistics in the high- tail of the distribution. APE then embeds the unmodified partonic hard scattering products in the event geometry, distributed according to the binary collision density as in Refs. [28, 12], computing trajectories for a fixed grid of initial azimuthal angles for the scattering axis at each production point. The partons are initially antiparallel by construction, as we include only the tree-level scattering in this work. This approximation does not affect single-inclusive distributions () but does affect acoplanarities; our report is thus the enhancement due to the impact of drift, which may be modified with a fluctuating treatment of the broadening.
Jet seeds in the plasma are evolved along a path parameterized by pathlength using the following kinematic workflow:
1.
Propagate the parton over a small timestep of length fm in the plasma center of mass rest frame.
2.
Compute the momentum transferred to the parton over the small pathlength due to enabled effects, updating parton momentum accordingly.
3.
Repeat until the parton escapes the plasma perimeter at freezeout.
We keep the quark masses, noting that for the light flavors considered here . For every initial condition, we perform separate simulations with contributions of radiative energy loss, collisional energy loss, and jet drift effects switched on or off, resulting in fully-correlated distributions of hard particles for each combination of effects. This minimizes Monte Carlo noise in relative comparisons between them, allowing extraction of statistics-hungry measures like the elliptic flow enhancement due to drift, (Fig. 3) in smaller sample sizes. Initial- and final- state properties of the hard partons are recorded before applying hadronization. We consider no medium response effects at this time.
Neglecting the running of , we tune the coupling for the two energy loss cases (both including drift) via reduced fit to CMS data at 0-5%centrality and [29]. To two significant figures, we find with radiative energy loss alone and with radiative and collisional energy loss, consistent with values from the literature [30]. We have also explored the use of various forms of cold nuclear matter effects, but the observables reported here were found to be largely insensitive to their inclusion. Herein we choose to follow the perturbative treatment of Refs. [2, 31, 32] for the Cronin effect, isospin effects, and nuclear shadowing, but we neglect cold nuclear matter energy loss.
A perturbative calculation of parton drift and energy loss requires input for the mean free path and Debye mass . The Debye screening mass is where we use , and by definition , with () the total cross section for the hard parton to interact with a quark (gluon) in the plasma and () the density of quarks (gluons) in the medium. Here, is chosen to be the total Gyulassy-Wang cross section , modeled after static scattering centers in the medium [33], and the cross sections and partial quark and gluon densities are taken from [34], which approximates the medium as an ideal gas of gluons and quarks. It should be noted that the identified parton number densities are not provided by the hydrodynamic equation of state and are necessary to compute the various flavor-dependent channels of jet-medium interactions. We set when the medium cools below the hadronization temperature, which is chosen to be MeV.
Calculations of the energy loss of high-energy partons due to stimulated gluon emission are computationally demanding. This has led to a range of approaches to evaluate energy loss in realistic media. Many of the traditional works select an analytic form for the plasma [35]; others create effective lengths and temperatures as realistic envelopes on an analytic model [36]; while others simplify or parameterize the energy loss form [3]. We elect to numerically compute single-emission Gyulassy-Levai-Vitev radiative energy loss at first order in opacity, including finite kinematic bounds for the momentum transfer to the medium and the radiated gluon momentum, using the Gyulassy-Wang form for the scattering potentials [35] at the level of the average. The behavior of the radiative energy loss is qualitatively different at low for different choices of kinematic bounds, where sub-eikonal corrections become large. In a medium of effective length , the mean energy loss rate is
with , , and [37, 35]. Here, is the quadratic Casimir of the representation of SU(3) ( for gluons, for quarks). We take fm for the thermalization time and do not treat the short-pathlength corrections as derived in Ref. [38], which may be significant even for large collision systems. Thus, for massless partons, , with the time relative to the collision in the plasma center of mass rest frame.
As described in Eq. (II), we evaluate the mean energy loss of a hard parton of energy in a QGP of fixed length and temperature, then compute the rate in by central differences. This yields the tabulated instantaneous energy loss rate we sample at runtime in each timestep of the partonic propagation. This method preserves the quadratic pathlength dependence associated with the Landau-Pomeranchuk-Migdal effect [39, 40] and the additional effects associated with the expanding medium, locally samples the properties of the QGP at each point of the partonic trajectory in a hydrodynamic geometry, and is computationally cheap at runtime.
We implement an analytic calculation for the collisional energy loss [41] and can include it or neglect it in the computation, so as to estimate the effect on the phenomenology of drift:
where is the thermal gluon mass. A finite-temperature perturbative calculation [42] shows only minor corrections to Eq. (4). Likewise, the additional finite- corrections to Ref. [42] were also shown to be small [43].
We treat the flow-mediated drift at the level of the average computed in Ref. [11]. The momentum transferred to the parton transverse to its trajectory is
Here, is the mean free path, is the Debye screening mass, and is the collective flow velocity, with components and to the parton momentum vector, and is the corresponding unit vector.
The distribution of leading hadrons can be calculated from the distribution of leading partons by convolution with fragmentation functions (FFs) [44]
In order to evaluate this expression for a discrete distribution of quarks and gluons, we perform an event-by-event fragmentation such that for a given parton with fixed flavor and , we sample from a probability distribution function for hadronic momentum fraction proportional to the inclusive fragmentation function, . We compute the distribution of inclusive charged hadrons, applying the JAM Collaboration’s “JAM20-SIDIS” FF fits for charged hadrons [45] accessed through LHAPDF [46].
This treatment of the system as a whole takes the most conservative choices for each phase of the computation. The results represent a realistic lower-bound on the effect size of flow mediated drift. This energy regime, and especially the flow-mediated drift effect, is quite sensitive to model choices at every step. To study the impact of various other viable scenarios, we also implement a sharper, more anisotropic optical Glauber initial condition for the medium and formulate a simple treatment of hybrid coalescence effects. For the initial entropy density of the sharp optical Glauber medium model, we take the harmonic mean of two Woods-Saxon nuclear density profiles [21, 22], . The higher anisotropy and sharper gradients of these initial conditions produce large deflections due to flow. The medium evolution is treated identically. Following [47], we presume that pions dominate the charged hadron distribution, thus we take the distribution of hadrons formed from the hybrid coalescence of a hard parton and a thermal parton to be
where we have taken the same model form for the Wigner distribution of the pion as [47] (with the momentum width of the pion GeV, the volume of the QGP, and the formation probability). Our thermal parton distribution is a simple Boltzmann distribution with the appropriate boson or fermion statistics for the hard parton’s color neutralizing counterpart,
III Results.
A simultaneous description of the suppression of jets and their elliptic modulation, , has been one of the benchmark problems of jet physics in heavy ion collisions [3, 4, 5, 6, 7, 8, 48]. Various microscopic models and calculations can simultaneously capture both observables in the high-energy regime, GeV, but perturbative calculations have not successfully described the intermediate and low energy regimes, GeV, where elliptic flow is generally underpredicted [4, 6].
Jet drift is a perturbative correction to the transverse momentum distribution that naturally supplies some of the missing flow. While this sub-eikonal effect does not directly modify the energy loss, drift may indirectly affect the suppression by deflecting the paths toward regions of lower density. This deflection, however, directly modifies the azimuthal distribution of particles. The coupling of drift to the medium correlates leading hadrons to the event plane by dragging hard particles towards “attractors” at the minor axes of the elliptic medium geometry, as schematically shown in Fig 1.
We report the suppression as computed from Eq. (1) in Fig 2 and the elliptic flow computed from Eq. (2) in Fig 3. We take the denominator of Eq. (1) as the number distribution of partons seeded from Pythia and compute by importance sampling as a discrete sum over all trajectories producing final state particles in the given bin.
where is the weight associated with parton and the final azimuthal angle of the trajectory relative to the computed hard event plane angle at each , within each event. Finally, the angle brackets represent an average over events.
Intriguingly, we obtain a nuclear modification factor in Fig. 2 which is unmodified due to the effects of drift. This suggests that any systematic bias of trajectories toward lower density is not significant for the nuclear modification factor. We do, however, see the anticipated enhancement of the elliptic flow shown in Fig. 3, which turns out to be sizeable: an enhancement of several times the experimental uncertainty (e.g. for charged pion at 3.00 - 3.25 GeV in the 40-50% bin [49]). Interestingly, the enhancement from drift qualitatively matches the excess in elliptic flow seen in experiment below 10 GeV [50]. While we make no claim that jet drift is the solution to the puzzle, it must nevertheless be an important part of the solution in the GeV regime.
We also note that the jet drift effect is strongest in semi-peripheral collisions (30-50% centrality), which balance a larger event eccentricity with the decreasing event temperatures in smaller-sized QGP. This demonstrates the interplay between energy loss and drift, whereby energy loss enhances the strength of drift, but strong energy loss extinguishes partons that have experienced strong drift. Likewise, the addition of collisional energy loss (and the refitting of the coupling) reduces the strength of the enhancement by about 50%, as shown in the ratio inset of Fig. 3, as a lower overall coupling is necessary to fit high- suppression, and the relative strength of drift is reduced. The addition of other sources of energy loss (including from cold nuclear matter effects) similarly reduces the relative strength of drift.
The strength of the anisotopic flow modification is significantly enhanced by the inclusion of coalescence or the use of a sharper initial condition for the medium evolution. As shown in Fig. 4, coalescence upshifts the of strongly drifted partons, expanding the region of interest and dramatically increasing the magnitude of the enhancement. This implementation takes an isotropic distribution of thermal partons, and thus does not include direct convolution with the azimuthal anisotropy of the medium, which is expected to have additional complex interplay with drift. The combination of both changes represents an optimistic, rational estimate, but is by no means an upper bound on the effect.
Likewise, the correlated deflections of particles from drifting in the same event geometry results in an enhancement of the acoplanarity (a deflection away from back-to-back) (Fig. 5) between the hadrons. The additional contribution from drift can be easily understood from Fig. 1, in which widely-separated parton trajectories are deflected toward the same “attractor” in the event geometry. While measures deflections correlated with the event plane, the azimuthally averaged acoplanarity provides a complementary measure of the magnitude of the deflections, independent of the event plane, although the strength of the enhancement varies azimuthally relative to the event plane.
We study acoplanarities in APE by embedding the pairs of partons from each hard process on exactly antiparallel trajectories at the same production point in a given event geometry. The resulting deflections of single partons, defined as the difference in initial and final trajectory angle, and acoplanarities, defined as the smallest angle between parton pairs from the same hard process in the final state, are plotted in Fig. 5. Like the elliptic flow , the mean deflection angles and acoplanarities show sizeable enhancements due to drift at low . Importantly, however, the centrality dependence has changed between Fig. 3 and Fig. 5. The elliptic flow enhancement is largest for the intermediate centrality bin, while the deflections and acoplanarities are largest in the central bin. This is a reflection of the fact that acoplanarities are sensitive to the magnitude of drift, which is largest in central collisions but is much more weakly correlated with the event plane. The elliptic flow, on the other hand, reflects not only the temperature but also the ellipticity, which is largest in peripheral collisions.
IV Conclusions.
Flow-mediated drift, as studied here, belongs both to a family of sub-eikonal effects that extend the theory of jet quenching toward lower , and to a family of antisymmetric effects that can generate azimuthal anisotropies of hard probes. This work shows that anisotropic effects modify hard observables in ways that survive event averaging, and sub-eikonal effects can systematically extend perturbative calculations to intermediate- and low-. We have further demonstrated that inclusive measures like and correlations like acoplanarities exhibit different systematics in their response to drift. Jet drift and similar effects are sensitive to properties of the medium that neither isotropic measures of hard probes nor soft-sector anisotropies can capture, including local flow structure, flowing hot spots, and other internal structural differences. These novel properties of jet drift provide an additional dimension of sensitivity with which to expand the prospects of jet QGP tomography championed by the field.
Throughout this study, the main results presented a conservative estimate of the effect of jet drift. We disabled in-medium interactions below the hadronization temperature, which removes the potential effect of late-time interactions where collective flow, and thus drift, is strongest. Despite these unfavorable assumptions, we show here that jet drift nonetheless leaves measurable imprints of the flow on low- particles, systematically enhancing elliptic flow and increasing dijet acoplanarity. The reasonable inclusion of coalescence and use of a sharper medium model significantly enhances the effect and serves to emphasize how sensitive the results of an intermediate simulation are to model choices, especially in regards to drift. Consequently, this study constitutes the first proof-of-principle demonstration of velocity tomography in event-by-event heavy-ion collisions and the first comprehensive phenomenological investigation of the flow modification of hard probes.
In the future, we plan to study the implications of drift for intra-jet particle and energy distributions. Previous work has considered that gradients and flow may impact the distribution of radiation in the jet cone [19, 11, 51]. Jet substructure is expected to be even more sensitive to flow than the inclusive observables considered here due to the combination of asymmetric gluon radiation and drift.
V Acknowledgments.
Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Award Number DE-SC0024560 (JB, HR, MS, IV) and the Los Alamos National Laboratory (Contract No. 89233218CNA000001) and its LDRD Program under Award Number 20240131ER (IV). This work utilized resources from the New Mexico State University High Performance Computing Group, which is directly supported by the National Science Foundation (OAC-2019000), the Student Technology Advisory Committee, and New Mexico State University and benefits from inclusion in various grants (DoD ARO-W911NF1810454; NSF EPSCoR OIA-1757207; Partnership for the Advancement of Cancer Research, supported in part by NCI grants U54 CA132383 (NMSU)). This research utilized services provided by the OSG Consortium [52, 53, 54, 55], which is supported by the National Science Foundation awards #2030508 and #1836650. The open-source APE software suite and associated resources are publicly available on GitHub. JB and IV acknowledge INT support.References
- [1] M. Connors, C. Nattrass, R. Reed & S. Salur, “Jet measurements in heavy ion physics”, Rev. Mod. Phys. 90, 025005 (2018), arXiv:1705.01974.
- [2] I. Vitev & M. Gyulassy, “High tomography of + Au and Au+Au at SPS, RHIC, and LHC”, Phys. Rev. Lett. 89, 252301 (2002), hep-ph/0209161.
- [3] J. Noronha-Hostler, B. Betz, J. Noronha & M. Gyulassy, “Event-by-event hydrodynamics jet energy loss: A solution to the puzzle”, Phys. Rev. Lett. 116, 252301 (2016), arXiv:1602.03788.
- [4] D. Molnar & D. Sun, “High-pT suppression and elliptic flow from radiative energy loss with realistic bulk medium expansion”, arXiv:1305.1046.
- [5] W. Zhao, W. Ke, W. Chen, T. Luo & X.-N. Wang, “From Hydrodynamics to Jet Quenching, Coalescence, and Hadron Cascade: A Coupled Approach to Solving the RAAv2 Puzzle”, Phys. Rev. Lett. 128, 022302 (2022), arXiv:2103.14657.
- [6] X. Zhang & J. Liao, “Jet Quenching and Its Azimuthal Anisotropy in AA and possibly High Multiplicity pA and dA Collisions”, arXiv:1311.5463.
- [7] B. Z. Kopeliovich, J. Nemchik, I. K. Potashnikova & I. Schmidt, “Quenching of high-pT hadrons: Energy Loss vs Color Transparency”, Phys. Rev. C 86, 054904 (2012), arXiv:1208.4951.
- [8] C. Andres, N. Armesto, H. Niemi, R. Paatelainen & C. A. Salgado, “Jet quenching as a probe of the initial stages in heavy-ion collisions”, Phys. Lett. B 803, 135318 (2020), arXiv:1902.03231.
- [9] CMS Collaboration, A. Hayrapetyan et al., “Overview of high-density QCD studies with the CMS experiment at the LHC”, arXiv:2405.10785.
- [10] ALICE Collaboration, S. Acharya et al., “Observation of Medium-Induced Yield Enhancement and Acoplanarity Broadening of Low-pT Jets from Measurements in pp and Central Pb-Pb Collisions at sNN=5.02 TeV”, Phys. Rev. Lett. 133, 022301 (2024), arXiv:2308.16131.
- [11] A. V. Sadofyev, M. D. Sievert & I. Vitev, “Ab Initio Coupling of Jets to Collective Flow in the Opacity Expansion Approach”, arXiv:2104.09513.
- [12] L. Antiporda, J. Bahder, H. Rahman & M. D. Sievert, “Jet Drift and Collective Flow in Heavy-Ion Collisions”, arXiv:2110.03590.
- [13] X.-N. Wang & M. Gyulassy, “HIJING: A Monte Carlo model for multiple jet production in p p, p A and A A collisions”, Phys. Rev. D 44, 3501 (1991).
- [14] I. P. Lokhtin, L. V. Malinina, S. V. Petrushanko, A. M. Snigirev, I. Arsene & K. Tywoniuk, “Heavy ion event generator HYDJET++ (HYDrodynamics plus JETs)”, Comput. Phys. Commun. 180, 779 (2009), arXiv:0809.2708.
- [15] K. Zapp, J. Stachel & U. A. Wiedemann, “JEWEL - a Monte Carlo Model for Jet Quenching”, PoS High-pT physics09, 022 (2009), arXiv:0904.4885.
- [16] J. H. Putschke et al., “The JETSCAPE framework”, arXiv:1903.07706.
- [17] D. Zigic, I. Salom, J. Auvinen, P. Huovinen & M. Djordjevic, “DREENA-A framework as a QGP tomography tool”, Front. in Phys. 10, 957019 (2022), arXiv:2110.01544.
- [18] J. Casalderrey-Solana, D. C. Gulhan, J. G. Milhano, D. Pablos & K. Rajagopal, “A Hybrid Strong/Weak Coupling Approach to Jet Quenching”, JHEP 1410, 019 (2014), arXiv:1405.3864, [Erratum: JHEP 09, 175 (2015)].
- [19] N. Armesto, C. A. Salgado & U. A. Wiedemann, “Measuring the collective flow with jets”, Phys. Rev. Lett. 93, 242301 (2004), hep-ph/0405301.
- [20] J. S. Moreland, J. E. Bernhard & S. A. Bass, “Bayesian calibration of a hybrid nuclear collision model using p-Pb and Pb-Pb data at energies available at the CERN Large Hadron Collider”, Phys. Rev. C 101, 024911 (2020), arXiv:1808.02106.
- [21] J. S. Moreland, J. E. Bernhard & S. A. Bass, “Alternative ansatz to wounded nucleon and binary collision scaling in high-energy nuclear collisions”, Phys. Rev. C92, 011901 (2015), arXiv:1412.4708.
- [22] J. S. Moreland, J. E. Bernhard, W. Ke & S. A. Bass, “Flow in small and large quark-gluon plasma droplets: the role of nucleon substructure”, Nucl. Phys. A 967, 361 (2017), arXiv:1704.04486.
- [23] H. Song, S. A. Bass & U. Heinz, “Viscous QCD matter in a hybrid hydrodynamic+Boltzmann approach”, Phys. Rev. C 83, 024912 (2011), arXiv:1012.0555.
- [24] S. A. Bass et al., “Microscopic models for ultrarelativistic heavy ion collisions”, Prog. Part. Nucl. Phys. 41, 255 (1998), nucl-th/9803035.
- [25] When discussing the power counting of the eikonal expansion, we refer to the parton energy rather than to avoid confusion with drift transverse to the jet axis. For light particle production at mid rapidity , the two are equivalent: .
- [26] J. Noronha-Hostler.
- [27] C. Bierlich et al., “A comprehensive guide to the physics and usage of PYTHIA 8.3”, arXiv:2203.11601.
- [28] J. E. Bernhard, J. S. Moreland, S. A. Bass, J. Liu & U. Heinz, “Applying Bayesian parameter estimation to relativistic heavy-ion collisions: simultaneous characterization of the initial state and quark-gluon plasma medium”, Phys. Rev. C94, 024907 (2016), arXiv:1605.03954.
- [29] CMS Collaboration, V. Khachatryan et al., “Charged-particle nuclear modification factors in PbPb and pPb collisions at TeV”, JHEP 1704, 039 (2017), arXiv:1611.01664.
- [30] Y. He, I. Vitev & B.-W. Zhang, “ Analysis of Inclusive Jet and di-Jet Production in Heavy Ion Reactions at the Large Hadron Collider”, Phys. Lett. B 713, 224 (2012), arXiv:1105.2566.
- [31] M. Gyulassy, P. Levai & I. Vitev, “Reaction operator approach to multiple elastic scatterings”, Phys. Rev. D 66, 014005 (2002), nucl-th/0201078.
- [32] J.-w. Qiu & I. Vitev, “Resummed QCD power corrections to nuclear shadowing”, Phys. Rev. Lett. 93, 262301 (2004), hep-ph/0309094.
- [33] M. Gyulassy & X.-n. Wang, “Multiple collisions and induced gluon Bremsstrahlung in QCD”, Nucl. Phys. B420, 583 (1994), nucl-th/9306003.
- [34] M. D. Sievert, I. Vitev & B. Yoon, “A complete set of in-medium splitting functions to any order in opacity”, Phys. Lett. B 795, 502 (2019), arXiv:1903.06170.
- [35] M. Gyulassy, I. Vitev & X. N. Wang, “High p(T) azimuthal asymmetry in noncentral A+A at RHIC”, Phys. Rev. Lett. 86, 2537 (2001), nucl-th/0012092.
- [36] C. Faraday & W. Horowitz, “Energy Loss in Small Collision Systems”, PoS HardProbes2023, 131 (2024), arXiv:2307.08355.
- [37] M. Gyulassy, P. Levai & I. Vitev, “Reaction operator approach to nonAbelian energy loss”, Nucl.Phys. B594, 371 (2001), nucl-th/0006010.
- [38] I. Kolbe & W. A. Horowitz, “Short path length corrections to Djordjevic-Gyulassy-Levai-Vitev energy loss”, Phys. Rev. C 100, 024913 (2019), arXiv:1511.09313.
- [39] A. B. Migdal, “Bremsstrahlung and pair production in condensed media at high-energies”, Phys. Rev. 103, 1811 (1956).
- [40] L. D. Landau & I. Pomeranchuk, “Limits of applicability of the theory of bremsstrahlung electrons and pair production at high-energies”, Dokl. Akad. Nauk Ser. Fiz. 92, 535 (1953).
- [41] E. Braaten & M. H. Thoma, “Energy loss of a heavy quark in the quark - gluon plasma”, Phys. Rev. D 44, R2625 (1991).
- [42] M. H. Thoma & M. Gyulassy, “Quark damping and energy loss in the high temperature QCD”, Nuclear Physics B 351, 491 (1991).
- [43] M. Djordjevic, “Collisional energy loss in a finite size QCD matter”, Phys. Rev. C 74, 064907 (2006), nucl-th/0603066.
- [44] C. Markert, R. Bellwied & I. Vitev, “Formation and decay of hadronic resonances in the QGP”, Phys. Lett. B 669, 92 (2008), arXiv:0807.1509.
- [45] Jefferson Lab Angular Momentum (JAM) Collaboration, E. Moffat, W. Melnitchouk, T. C. Rogers & N. Sato, “Simultaneous Monte Carlo analysis of parton densities and fragmentation functions”, Phys. Rev. D 104, 016015 (2021), arXiv:2101.04664.
- [46] A. Buckley, J. Ferrando, S. Lloyd, K. Nordström, B. Page, M. Rüfenacht, M. Schönherr & G. Watt, “LHAPDF6: parton density access in the LHC precision era”, Eur. Phys. J. C 75, 132 (2015), arXiv:1412.7420.
- [47] V. Greco, C. M. Ko & P. Levai, “Parton coalescence and anti-proton / pion anomaly at RHIC”, Phys. Rev. Lett. 90, 202302 (2003), nucl-th/0301093.
- [48] L. S. Moriggi, E. S. Rocha & M. V. T. Machado, “Study of the azimuthal asymmetry in heavy ion collisions combining initial state momentum orientation and final state collective effects”, Phys. Rev. D 108, 074013 (2023).
- [49] ALICE Collaboration, S. Acharya et al., “Anisotropic flow of identified particles in Pb-Pb collisions at TeV”, JHEP 1809, 006 (2018), arXiv:1805.04390.
- [50] CMS Collaboration, A. M. Sirunyan et al., “Azimuthal anisotropy of charged particles with transverse momentum up to 100 GeV/ c in PbPb collisions at =5.02 TeV”, Phys. Lett. B 776, 195 (2018), arXiv:1702.00630.
- [51] J. a. Barata, J. G. Milhano & A. V. Sadofyev, “Picturing QCD jets in anisotropic matter: from jet shapes to energy energy correlators”, Eur. Phys. J. C 84, 174 (2024), arXiv:2308.01294.
- [52] R. Pordes, D. Petravick, B. Kramer, D. Olson, M. Livny, A. Roy, P. Avery, K. Blackburn, T. Wenaus, F. Würthwein, I. Foster, R. Gardner, M. Wilde, A. Blatecky, J. McGee & R. Quick, “The open science grid”, in “J. Phys. Conf. Ser.”, p. 012057.
- [53] I. Sfiligoi, D. C. Bradley, B. Holzman, P. Mhashilkar, S. Padhi & F. Wurthwein, “The pilot way to grid resources using glideinWMS”, in “2009 WRI World Congress on Computer Science and Information Engineering”, p. 428–432.
- [54] OSG, “OSPool”.
- [55] OSG, “Open Science Data Federation”, https://osdf.osg-htc.org/.